OpenStudy (anonymous):

In the equation ax + cy = n -bx + ry = n Solve by the following process If you eliminate x what will we have to multiply both equations by!

OpenStudy (anonymous):

first one by b, second one by a

OpenStudy (anonymous):

get \[abx+bcy=bn\] \[-abx+ary=an\]

OpenStudy (anonymous):

add and the x's will drop out \[(bc+ar)y=(a+b)n\]

OpenStudy (anonymous):

\[y=\frac{(a+b)n}{bc+ar}\]

OpenStudy (anonymous):

Could you show step by step?

OpenStudy (anonymous):

well actually those are the steps. you want the x's to drop out. the first equation shows ax and the second shows -bx you want the coefficients to work out so that one is the negative of the other

OpenStudy (anonymous):

if you multiply the first one by b you bet abx and if you multiply the second one by a you get -abx and now one is the negative of the other

OpenStudy (anonymous):

for example if one had 5x and the other had -3x you would multiply the top one by 3 and the bottom one by 5 to get 15x and -15x then they would "cancel' (add up to zero) when you added the equations

OpenStudy (anonymous):

ax + cy = n -bx + ry = n multiply first by B and second by a?